In a local theatre, 2000 tickets can be sold in a week when the price of a movie is set at $10 week. According to market research, Ed (price elasticity of demand) is 0.8. Now the management of the theatre has decided to raise the price of a movie to be $12.

(The formula of Ed is for your reference.)

37) The weekly demand of tickets is going to __________.

A)             decrease by 10%

B)             decrease by 12%

C)      decrease by 14%

D)      decrease by 16%

E)      decrease by 18%

38) The weekly revenue is going to __________.

A)             decrease by $260

B)             have no change

C)      increase by $160

D)      increase by $260

E)      increase by $360
Answer #1

1.

D) decrease by 16%

Explanation :

Price elasticity = Percentage change in quantity demanded /percentage change in price

0.8=percentage change in quantity demanded /20

16=percentage change in quantity demanded.

Percentage change in price =(P2 - P1) /P1

=(12-10)/12]

=2/10

=20%

2.

C) increase by $160

Explanation :

Total revenue when price is 10.

TR =Q *P

=10*2000

=20,000

Total revenue when price is 12.

Quantity demanded will decrease by 16%.

2000-16%=1680.

TR =Q*P

=1680*12

=20160.

So total revenue is increase by=20,160-20,000=160.